If (2,1,-3), (-2,3,4) & (1,2,2) are the midpoints of the sides
BC, CA, AB of a triangle ABC, then the coordinates of it’s vertex A are

Grade:12th Pass

1 Answers

Vikas TU
6 years ago
Dear Student,
Mid point rule gives (A + B)/2 = (1,2,2) , (B + C)/2
= (2,1,-3) amd (C + A)/2 = (-2,3,4).
Solving the equations we get A = (-3,4,9).
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)